OpenStudy (rainbowdash1145):

What is the Molecular Mass of the following? SO2 O3 CO2 CH4 N2

OpenStudy (rainbowdash1145):

The top one should actually be SO3

OpenStudy (rainbowdash1145):

@phi @.Sam. @AloneS

OpenStudy (rainbowdash1145):

The only one i can think is O3=43 grams/mole

OpenStudy (s4sensitiveandshy):

First of all write down the atomic mass of each element. \[CO_2 = Atomic ~mass~of~C + (total) atomic ~mass~of~ O\] C= 12.01 amu O=16 because there are two elements of O, multiply 16 by 2 12.01+(16 x 2) = answer

OpenStudy (s4sensitiveandshy):

\[O_3= O+O+O = O *3 \]

OpenStudy (rainbowdash1145):

Thank You... This really helps.

OpenStudy (rainbowdash1145):

So CO2 is 44.01?

OpenStudy (rainbowdash1145):

@S4Sensitiveandshy

OpenStudy (s4sensitiveandshy):

yes sir.

OpenStudy (rainbowdash1145):

Thanks again!

OpenStudy (s4sensitiveandshy):

Good job. you're welcome.